Please help me solve this word problem. The formula for the volume of a cone was V= 1/3πr²h. If Andrew can jam 62.8 cubic inches of ice cream into a waffle cone (with none heaping over) and cone's diameter is 8 inches, how tall is the cone?
Answer by stanbon(75887)

Please help me solve this word problem. The formula for the volume of a cone was V= 1/3πr²h. If Andrew can jam 62.8 cubic inches of ice cream into a waffle cone (with none heaping over) and cone's diameter is 8 inches, how tall is the cone?
----------------
62.8 = (1/3)pi*4^2 * h
------
16pih = 3*62.8
----
h = 188.4/50.27
---
height = 3.75 inches
-------------
Cheers,
Stan H.
-------------